40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 2392     Entry time: Thu Dec 10 18:27:48 2009
Author: Mott 
Type: Update 
Category: General 
Subject: Updated R & T Measurements 

Attached are updated plots of the T&R Measurements for a variety of mirrors, and diagrams for the setup used to make the measurements.

T is plotted for the 1064 nm measurement, since these mirrors are highly reflective at 1064, and either R or R&T are plotted for the 532 nm measurement, depending on how larger the R signal is.

As with the previous set of plots, the error bars here are purely statistical, and there are certainly other sources of error not accounted for in these plots.  In general, the T measurement was quite stable, and the additional errors
are probably not enormous, perhaps a few percent.

The mirrors are:

Y1-1037-00.50CC

Y1-2037-45S

Y1-2037-45P

Y1S-1025-0

Y1S-1025-45

 

Attachment 1: Y1S-0.png  7 kB  | Hide | Hide all
Y1S-0.png
Attachment 2: Y1S-45.png  8 kB  | Hide | Hide all
Y1S-45.png
Attachment 3: Y45P.png  5 kB  | Hide | Hide all
Y45P.png
Attachment 4: Y45S.png  6 kB  | Hide | Hide all
Y45S.png
Attachment 5: Y150CC.png  6 kB  | Hide | Hide all
Y150CC.png
Attachment 6: Setup.png  41 kB  | Hide | Hide all
Setup.png
ELOG V3.1.3-